OpenStudy (anonymous):

can someone check this for me? 1. Evaluate: −7 − 8 (4 points) –1 –15 1 15 2. Evaluate: −6 − (−4) (4 points) –10 –2 24 2 3. Evaluate: −10 + 16 (4 points) –4 6 26 –6 4. Evaluate: 3^2 + (6 − 2) · 4 − 6/3 (4 points) 20 23 50 26

OpenStudy (anonymous):

1.B 2.D 3.B 4.C

jimthompson5910 (jim_thompson5910):

# 1 is correct since -7 - 8 = -15

jimthompson5910 (jim_thompson5910):

#2 is not though

jimthompson5910 (jim_thompson5910):

-6 - (-4) -6 + 4 -2 so -6 - (-4) = -2

OpenStudy (anonymous):

Thank you :)

jimthompson5910 (jim_thompson5910):

# 3 is correct since −10 + 16 = 16 - 10 = 6

jimthompson5910 (jim_thompson5910):

but #4 is not 3^2 + (6 − 2) · 4 − 6/3 3^2 + (4) · 4 − 6/3 9 + (4) · 4 − 6/3 9 + 16 − 6/3 9 + 16 − 2 25 − 2 23 So 3^2 + (6 − 2) · 4 − 6/3 = 23

OpenStudy (anonymous):

Thank you. Can you check 3 more?

OpenStudy (anonymous):

5. Evaluate: 3 · 4 − 8 ÷ 2 (4 points) 2 8 3 0 6. Evaluate: 4 − 28 ÷ 7 + 3^2 (4 points) –9 9 5.6 6 7. Evaluate: ([30 + 6] − 3^2) ÷ 9 · 2 (4 points) 6.6 6 1.5 34

jimthompson5910 (jim_thompson5910):

which answers did you get for these

OpenStudy (anonymous):

5.B 6.C. 7.B

jimthompson5910 (jim_thompson5910):

# 5 is correct

jimthompson5910 (jim_thompson5910):

# 6 is not

jimthompson5910 (jim_thompson5910):

4 − 28 ÷ 7 + 3^2 4 − 28 ÷ 7 + 9 4 − 4 + 9 0 + 9 9 So 4 − 28 ÷ 7 + 3^2 = 9

jimthompson5910 (jim_thompson5910):

# 7 is correct

OpenStudy (anonymous):

Thank you so much :) I will be posting more math questions I have another open math question right now if you want to come help me : )

jimthompson5910 (jim_thompson5910):

alright, I'll check it out

OpenStudy (anonymous):

Thanks again :)

jimthompson5910 (jim_thompson5910):

yw